I don't get linear equations, slopes, rate. Or whatever x = y, or m=, b=. Like these problems: 3x-y=8.
I don't know how to do them, I'm in Pre-Alegbra. I want to be in Algebra 1, but my teacher is Asian! I can not understand her, therefore, I do not know how to do this. Please, help! I really want to pass 8th grade, and be in Algebra 1.

Answer by richard1234(7193)

The nationality of your teacher shouldn't affect your ability to learn; you'll have to learn to cope with it. The high school I go to has teachers that are mostly Turkish or Russian.

The equation "y = mx + b" is simply a *general* form of a line. For example, 3x - y = 8 becomes y = 3x - 8 (after some algebraic manipulation), in which m = 3 (i.e. the slope is 3) and the y-intercept is -8 (i.e. the graph crosses the y-axis at (0,-8)).